A rectangle with height and width equal to 4 and 25 respectively, is drawn on a checkered
paper. Bazil paints a random horizontal 1 x 3 rectangle, and Peter paints a random vertical 3 x 1
rectangle (each rectangle consists of 3 sells). Find the probability that at least one of the cells is
painted twice. Express the answer in percent, and round to the nearest integer.
1
Expert's answer
2020-11-17T16:39:48-0500

We assume that Bazil and Peter paint on the drawn rectangle. The picture below presents

rectangle "5\times6".



As we can see, it is possible to paint 20 different horizontal "1\times3" rectangles. In a similar way we conclude that on rectangle "4\times25" it is possible to paint "23*4=92" horizontal "1\times3" rectangles and "25*2=50" different verical "3\times1" rectangles. We have "4600" possible combinations in total. On "3\times3" square it is possible to paint 9 different intersections of vertical and horizontal rectangles. We can consider "23*2=46" different "3\times3" squares on the rectangle. Thus, we obtain "46*9=414" different combination, when two rectangles intersect each other. I.e., there will be cells painted twice. Thus, the probability that one cell is painted twice is:"p=\frac{414}{4600}=0.09".

Answer: 0.09